# On the Resilience of LLM-Based Multi-Agent Collaboration with Faulty Agents

### v3

@@ Appendix D Limitations
There are several limitations in this study. First, due to budget constraints, we explore only GPT-3.5 and GPT-4o. Since our primary goal is to fairly evaluate different multi-agent systems’ resilience against faulty agents, we believe the results would not greatly differ from other models. The second limitation is the selection of multi-agent systems and downstream tasks, which cannot be comprehensive. We mitigate this by selecting representative systems from three well-established human collaboration modes (Yang & Zhang, 2019; Alexy, 2022; Mihm et al., 2010) and using four commonly-used datasets for benchmarking the abilities of multi-agent systems (Liang et al., 2024; Chen et al., 2021). The final limitation concerns the analysis, where latent variables affecting system resilience might be unidentified. To minimize this risk, we examine system architectures, downstream tasks, error rates, error types, agent roles, and the number of agents’ communications. To the best of our knowledge, no additional factors influencing system resilience are found.
Appendix E Ethics Statements and Broader Impacts

### v4

@@ Appendix D Limitations
Our study offers the first systematic probe of multi-agent resilience, yet several caveats remain. (1) Due to budget constraints, all agents share either GPT-3.5 or GPT-4o back-bones with temperature zero, and evaluation relies exclusively on automated scores. Since our primary goal is to fairly evaluate different multi-agent systems’ resilience against faulty agents, we believe the results would not greatly differ from other models. (2) All experiments focus on four text-only benchmarks with automatic metrics, so our conclusions may not extrapolate to multimodal interaction, open-ended dialogue, or long-horizon planning. We mitigate this by selecting representative systems from three well-established human collaboration modes (Yang & Zhang, 2019; Alexy, 2022; Mihm et al., 2010) and using four commonly-used datasets for benchmarking the abilities of multi-agent systems (Liang et al., 2024; Chen et al., 2021). (3) Because there is no universal set of agent profiles that fits every structure, each system is run with the authors’ own code base and role design—e.g., a flat system has no leader, whereas the linear and hierarchical ones do. This inevitably entangles structural effects with prompt engineering. To mitigate bias, we (i) keep each framework exactly as released and tuned by its authors, and (ii) select two representative systems per structure so that variations in individual prompts partially average out. Nonetheless, residual prompt and implementation differences may still influence the measured resilience gaps.